# Gröbner Basis Cryptanalysis of Ciminion and Hydra

### v1

@@ Abstract
Ciminion and Hydra are two recently introduced symmetric key Pseudo-Random Functions for Multi-Party Computation applications. For efficiency both primitives utilize quadratic permutations at round level. Therefore, polynomial system solving-based attacks pose a serious threat to these primitives. For Ciminion we construct a quadratic degree reverse lexicographic (DRL) Gröbner basis for the iterated polynomial model via affine transformations. For Hydra we provide a computer-aided proof in SageMath that a quadratic DRL Gröbner basis is already contained within the iterated polynomial system for the Hydra heads after affine transformations and a linear change of coordinates.
Our Ciminion DRL Gröbner basis simplifies cryptanalysis, since one does not need to impose genericity assumptions, like being regular or semi-regular, anymore to derive complexity estimates on key recovery attacks.
In the Hydra proposal it was claimed that $r_{\mathcal{H}}=31$ rounds for the heads are sufficient to achieve $128$ bits of security against Gröbner basis attacks for key recovery. However, for $r_{\mathcal{H}}=31$ standard term order conversion to a lexicographic (LEX) Gröbner basis for our Hydra DRL Gröbner basis requires just $126$ bits. Moreover, via the Eigenvalue Method up to $r_{\mathcal{H}}=33$ rounds can be attacked below $128$ bits.
@@
Gröbner basis Eigenvalue Method Ciminion Hydra
@@ 1 Introduction
With secure Multi-Party Computation (MPC) several parties can jointly compute a function without revealing their private inputs. In practice, data often has to be securely transferred from and to third parties before it can be used in MPC protocols. Also, one may need to store intermediate results securely in a database. For efficiency, it is recommended to perform encryption via MPC-friendly Pseudo-Random Functions (PRF) [GRR+16] which utilize a secret-shared symmetric key. Moreover, one can utilize MPC-friendly PRFs to speed up MPC applications, e.g. [IKOS07, LTW13, CDG+17]. Compared to classical bit-based PRFs their MPC counterparts have two novel design criteria:
@@
Two such recently introduced PRFs are Ciminion [DGGK21] and Hydra [GØSW23]. These constructions are based on the iteration of degree $2$ Feistel and Lai–Massey permutations. Due to the low degree at round level, both designs consider Gröbner basis attacks [ACG+19] as one of the most threatening attack vectors.
Let us shortly review a Gröbner basis attack on an iterated design.
@@
    Model the function with an iterated system of polynomials
    $\mathcal{R}(\mathbf{x}_{i-1},\mathbf{y})-\mathbf{x}_{i}=\mathbf{0},$
    where $\mathcal{R}:\mathbb{F}_{q}^{n}\to\mathbb{F}_{q}^{n}$ is a round function. The $\mathbf{x}_{i}$’s are intermediate state variables with $\mathbf{x}_{0}$ the input state and $\mathbf{x}_{r}$ the output state of the function and $\mathbf{y}$ the cryptographic secret variables.
@@
    Convert to an elimination order, e.g. the lexicographic (LEX) term order.
@@
    Solve the univariate equation.
Let $I\subset P=\mathbb{F}_{q}[x_{1},\dots,x_{n}]$ be a zero-dimensional ideal, let $d=\dim_{\mathbb{F}_{q}}\left(P/I\right)$ denote the $\mathbb{F}_{q}$-vector space dimension of the quotient space, and assume that a DRL Gröbner basis for $I$ is known. The original FGLM algorithm [FGLM93] for term order conversion runs in $\mathcal{O}\left(n\cdot d^{3}\right)$. An improved probabilistic FGLM variant [FGHR14] runs in $\tilde{\mathcal{O}}\left(d^{\omega}\right)$, where $2.372862\leq\omega<$ 2.37286, if the ideal is in shape position, and in $\tilde{\mathcal{O}}\left(n\cdot d^{\omega}\right)$ if a LEX Gröbner basis of the radical ideal is computed. Further, a sparse linear algebra FGLM variant [FM17] achieves $\mathcal{O}\left(\frac{d^{2+\frac{n-1}{n}}}{\sqrt{n}}\right)$. To factor the univariate LEX polynomial efficiently Bariant et al. [BBLP22, §3.1] proposed to compute its GCD with the field equation. To improve from the generic GCD complexity they propose to perform an auxiliary division by remainder to reduce both polynomials to the same degree. Provided that just a view solutions stem from the base field $\mathbb{F}_{q}$, the complexity of $\mathbb{F}_{q}$-valued root extraction via the GCD is given by
$\mathcal{O}\Big{(}d\cdot\log(d)\cdot\log\big{(}\log(d)\big{)}\cdot\big{(}\log(d)+\log(q)\big{)}\Big{)},$ (1)
if $d\leq q$ else one has to replace the roles of $d$ and $q$ in the complexity estimate.
To estimate the complexity of Gröbner basis computations the Ciminion designers assumed that their polynomial systems are regular [Eis95, Chapter 17] and the Hydra designers assumed that their polynomial systems are semi-regular [Frö85, Par10]. Under these assumptions it is well-known that the complexity of Gröbner basis computations can be estimated via the degree of regularity [BFS04], but both designs lack a formal proof for their genericity assumptions.
Recently, Caminata & Gorla [CG21] introduced the notion of being in generic coordinates. Essentially, if a polynomial system is in generic coordinates, then the complexity of DRL Gröbner basis computations can always be expressed in terms of the Macaulay bound. We also note that for a regular polynomial system the degree of regularity is equal to the Macaulay bound.
In another recent work, Steiner [Ste24] analyzed several cryptographic polynomial systems of MiMC [AGR+16], Hades [GLR+20] and GMiMC [AGP+19] for being in generic coordinates. In particular, he derived an efficient criterion [Ste24, Theorem 3.2] to prove that a polynomial system is in generic coordinates. Moreover, in his analysis of GMiMC he revealed that iterated polynomial systems for Feistel polynomial systems can be verified to be in generic coordinates by computing the rank of an associated linear system. Our work can be seen as generalization of his work since Ciminion and Hydra essentially boil down to Feistel-like polynomial systems.
1.1 Our Results
For the iterated Ciminion polynomial system we prove in LABEL:Th:_Ciminion_Gr\"obner_basis that a DRL Gröbner basis can be constructed via affine transformations. This trivializes Step (2) in a Gröbner basis attack, and one can directly proceed with term order conversion or other polynomial system solving techniques to solve for the key.
It is also worthwhile mentioning that Bariant [Bar23] recently described that the univariate LEX polynomial of Ciminion can be constructed via univariate polynomial multiplication. We review this construction in more detail in Section 3.2, most importantly his attack can be prevented with additional key additions in the rounds of Ciminion. Moreover, if the keys are only added to the non-linear components, then we can maintain the iterated DRL Ciminion Gröbner basis (LABEL:Th:_Ciminion_Gr\"obner_basis). We call this slightly modified variant Ciminion2 (Section 3.3.1).
For Hydra, in Theorem 4.1 we derive a linear system to verify generic coordinates for the iterated Hydra polynomial system. In case the Hydra polynomial system is in generic coordinates, then we can also use the linear system to perform a change of coordinates and produce a quadratic polynomial system which contains a zero-dimensional DRL Gröbner basis. For Hydra this gives us to options to solve for the key.
  1. (I)
    We ignore the additional equations and just use the DRL Gröbner basis to proceed with polynomial system solving.
  2. (II)
    We exploit the quadratic DRL Gröbner basis to reduce the standard Macaulay matrix for the overdetermined Hydra polynomial system to a submatrix whose columns are only indexed by square-free monomials.
For (I), to compute the solutions of our polynomial systems we take a different approach than the classical term order conversion to LEX. We propose to compute the zeros via the eigenvalues of the so-called multiplication matrices, see Section 2.5. Then, the complexity of the Eigenvalue Method is dominated by the complexity of determinant computations. We also note that the probabilistic FGLM algorithm [FM17] boils down to matrix multiplications with the multiplication matrices, hence the generic Eigenvalue Method and the probabilistic FGLM algorithm share the same complexity estimate up to multiplicative constants. In particular, in Section 5 we discuss how the multiplicative structure of quadratic DRL Gröbner bases can be exploited to improve the complexity of the determinant for the characteristic polynomial. Consequently, this will yield the best currently known attacks on Hydra.
For (II), computing a DRL Gröbner basis for multivariate quadratic polynomial systems is an omnipresent problem in cryptography. However, for the full Hydra polynomial system we have an advantage over the general problem, because the system already contains a zero-dimensional DRL Gröbner basis. Therefore, to decrease the size of the Macaulay matrices in a Gröbner basis computation, we propose to perform an auxiliary division by remainder computation with respect to the Gröbner basis. Then, the remainders will only consist of square-free monomials, so we can reduce Gaussian elimination to a submatrix of the Macaulay matrix whose columns are only indexed by square-free monomials. We discuss this approach and its complexity in Section 6. We also note that the reduction to square-free Macaulay matrices for polynomial systems $\mathcal{F}\subset\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ was already discussed by Bardet et al. [BFSS13]. Over $\mathbb{F}_{2}$ one can always include the field equations $F=\left(x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\right)$ to restrict to the $\mathbb{F}_{2}$-valued solutions of $\mathcal{F}$. Obviously, the complexity of division by remainder with respect to $F$ is negligible compared to the cost of Gaussian elimination. This is not the case for our Hydra polynomial system, in particular a simple inequality reveals that for Hydra the auxiliary division by remainder step, i.e. the construction of square-free Macaulay matrices, dominates the complexity of DRL Gröbner basis computations (Equation 66). In addition, to highlight the analogy from polynomial system solving over $\mathbb{F}_{2}$, over an arbitrary field $K$ we call square-free Macaulay matrices $K$-Boolean Macaulay matrices.
We implemented Ciminion and Hydra polynomial systems in the computer algebra systems SageMath [Sag24] and OSCAR [OSC24]. [^1] https://github.com/sca-research/Groebner-Basis-Cryptanalysis-of-Ciminion-and-Hydra.git In particular, for Hydra we have implemented the linear system for the generic coordinates verification as well as the linear change of coordinates to produce a quadratic DRL Gröbner basis. We stress that our SageMath implementations is able to handle practical parameters like $128$ bit prime numbers and full rounds Hydra with $r_{\mathcal{H}}=39$.
Finally, in Section 7 we discuss how our results can be applied for cryptanalysis of Ciminion and Hydra.
  • For Ciminion2, utilizing the DRL Gröbner basis the cost of solving for the first key pair is proportional to
    $\mathcal{O}\left(d^{\omega\cdot(r_{C}+r_{E}-2)+1}\right).$ (2)
    Let $64\leq s\leq\log_{2}\left(q\right)$ denote the target security level of Ciminion, then one requires that $r_{E}\in\left\{s+6,\left\lceil\frac{2\cdot(s+6)}{3}\right\rceil\right\}$ [DGGK21, Table 1]. Since $\omega\geq 2$, this complexity will always exceed the target security level.
  • For Hydra, extracting a zero-dimensional DRL Gröbner basis and solving via the Eigenvalue Method turns out to be the most competitive attack vector. In [GØSW23, §7.4] the Hydra designers claim that $r_{\mathcal{H}}=31$ rounds are sufficient to achieve at least $128$ bits of security against Gröbner basis attacks. Note that this claim is already invalidated if we perform term order conversion to LEX via the probabilistic FGLM algorithm. Via the Eigenvalue Method we can attack up to $r_{\mathcal{H}}=33$ rounds below $128$ bit complexity. However, Hydra with full rounds $r_{\mathcal{H}}=39$ achieves $128$ bits security against all strategies analyzed in this paper. Our complexities for various attack strategies on Hydra are summarized in Table 1.
@@ 1.1 Our Results
Table 1: Summary of Hydra complexity estimations for various Gröbner basis attacks, see Tables 4, 5 and 6. All computations are performed over the prime $q=2^{127}+45$ with $\omega=2$.
@@
$r_{\mathcal{H}}$ Eigenvalue Method Term Order Conversion Proven Solving Degree Boolean Semi-Regular Estimate Semi-Regular Estimate [GØSW23]
$20$ $\mathbf{75}$ $82$ $105$ $91$ $89$
$29$ $\mathbf{111}$ $118$ $151$ $130$ $131$
$31$ $\mathbf{119}$ $126$ $161$ $140$ $142$
$33$ $\mathbf{127}$ $135$ $172$ $149$ $153$
$34$ $\mathbf{131}$ $139$ $177$ $154$ $159$
$39$ $\mathbf{151}$ $159$ $203$ $176$ $183$
@@ 2 Preliminaries
We denote fields with $K$ and their algebraic closure by $\bar{K}$. If the field and the number of variables are clear from context, then we abbreviate the polynomial ring as $P=K[x_{1},\dots,x_{n}]$. Matrices $\mathbf{M}\in k^{m\times n}$ are denoted with upper bold letters and vectors $\mathbf{v}\in\mathbb{F}_{q}^{n}$ by lower bold letters. Matrix-vector and matrix-matrix products are written as $\mathbf{M}\mathbf{v}$.
For polynomial ideals $I\subset K[x_{1},\dots,x_{n}]$, we denote the zero locus, i.e. its set of zeros, over $\bar{k}$ as
$\mathcal{Z}\left(I\right)=\left\{\mathbf{p}\in\bar{k}^{n}\mid f(\mathbf{p})=0,\ \forall f\in I\right\}\subset\mathbb{A}^{n}_{\bar{k}}.$ (3)
For homogeneous ideals $I$ we denote the projective zero locus over $\bar{K}$ by $\mathcal{Z}_{+}\left(I\right)\subset\mathbb{P}^{n-1}_{\bar{k}}$.
@@
Analog, we denote the homogenization of ideals $I^{\text{hom}}=\left\{f^{\text{hom}}\mid f\in I\right\}$ and of finite systems of polynomials $\mathcal{F}^{\text{hom}}=\left\{f_{1}^{\text{hom}},\dots,f_{m}^{\text{hom}}\right\}$.
For a homogeneous ideal $I\subset P$ and an integer $d\geq 0$ we denote
$I_{d}=\left\{f\in I\mid\deg\left(f\right)=d,\ f\text{ homogeneous}\right\}.$ (5)
For a term order $>$ and an ideal $I\subset P$ we denote with $\operatorname{in}_{>}(I)=\{\operatorname{LT}_{>}(f)\mid f\in I\}$ the initial ideal of $I$, i.e. the ideal of leading terms of $I$, with respect to $>$.
Every polynomial $f\in K[x_{1},\dots,x_{n}]$ of degree $d$ can be written as $f=f_{d}+f_{d-1}+\ldots+f_{0}$, where $f_{i}$ is homogeneous of degree $i$. We denote its highest degree component $f_{d}$ of $f$ with $f^{\text{top}}$, and analog we denote $\mathcal{F}^{\text{top}}=\left\{f_{1}^{\text{top}},\dots,f_{m}^{\text{top}}\right\}$. Note that
$f^{\text{top}}=f^{\text{hom}}\mod(x_{0}).$ (6)
Let $I,J\subset K[x_{1},\dots,x_{n}]$ be ideals, then we denote with
$I:J=\left\{f\in K[x_{1},\dots,x_{n}]\mid\forall g\in J\colon f\cdot g\in I\right\}$ (7)
the usual ideal quotient, and with
$I:J^{\infty}=\bigcup_{i\geq 1}I:J^{i}$ (8)
the saturation of $I$ with respect to $J$.
Let $I,\mathfrak{m}\in K[x_{0},\dots,x_{n}]$ be homogeneous ideals where $\mathfrak{m}=(x_{0},\dots,x_{n})$, then we call $I^{\text{sat}}=I:\mathfrak{m}^{\infty}$ the saturation of $I$.
Since we analyze vector-valued functions $F:\mathbb{F}_{q}^{n}\to\mathbb{F}_{q}^{n}$ we also denote vectors of polynomial ring variables as $\mathbf{x}=(x_{1},\dots,x_{n})^{\intercal}$. Moreover, for a term order $>$ on $P$, if not specified otherwise we always assume that the variables within a vector $\mathbf{x}$ are naturally ordered with respect to $>$, i.e. $x_{1}>\ldots>x_{n}$.
We will often encounter the lexicographic and the degree reverse lexicographic term order which we will abbreviate as LEX and DRL respectively.

### v4

@@ Abstract
Ciminion and Hydra are two recently introduced symmetric key Pseudo-Random Functions for Multi-Party Computation applications. For efficiency, both primitives utilize quadratic permutations at round level. Therefore, polynomial system solving-based attacks pose a serious threat to these primitives. For Ciminion, we construct a quadratic degree reverse lexicographic (DRL) Gröbner basis for the iterated polynomial model via linear transformations. With the Gröbner basis we can simplify cryptanalysis, as we no longer need to impose genericity assumptions to derive complexity estimates. For Hydra, with the help of a computer algebra program like SageMath we construct a DRL Gröbner basis for the iterated model via linear transformations and a linear change of coordinates. In the Hydra proposal it was claimed that $r_{\mathcal{H}}=31$ rounds are sufficient to provide $128$ bits of security against Gröbner basis attacks for an ideal adversary with $\omega=2$. However, via our Hydra Gröbner basis standard term order conversion to a lexicographic (LEX) Gröbner basis requires just $126$ bits with $\omega=2$. Moreover, using a dedicated polynomial system solving technique up to $r_{\mathcal{H}}=33$ rounds can be attacked below $128$ bits for an ideal adversary.
@@
Gröbner basis Ciminion Hydra
@@ 1 Introduction
With secure Multi-Party Computation (MPC) several parties can jointly compute a function without revealing their private inputs. In practice, data often has to be securely transferred from and to third parties before it can be used in MPC protocols. Also, one may need to store intermediate results securely in a database. For efficiency, it is recommended to perform encryption via an MPC-friendly Pseudo-Random Function (PRF) [GRR+16] which utilizes a secret-shared symmetric key. Compared to classical bit-based PRFs their MPC counterparts have two novel design criteria:
@@
Two such recently introduced MPC-friendly PRFs are Ciminion [DGGK21] and Hydra [GØSW23]. These constructions are based on the iteration of degree $2$ Feistel and Lai–Massey permutations. Due to the low degree at round level, both designs consider Gröbner basis attacks as one of the most threatening attack vectors.
In this paper we perform dedicated Gröbner basis cryptanalysis for both designs. Let us quickly sketch the main steps of a Gröbner basis attack on a symmetric design:
@@
    Model the function with a system of polynomials. For Ciminion and Hydra we analyze an iterated polynomial model $\mathcal{R}(\mathbf{x}_{i-1},\mathbf{y})-\mathbf{x}_{i}=\mathbf{0}$, where $\mathcal{R}:\mathbb{F}_{q}^{n}\to\mathbb{F}_{q}^{n}$ is the round function, the $\mathbf{x}_{i}$’s are intermediate state variables with $\mathbf{x}_{0}$ the input state and $\mathbf{x}_{r}$ the output state of the function, and $\mathbf{y}$ the cryptographic secret variables.
@@
    Construct a univariate polynomial, e.g. via term order conversion to the lexicographic (LEX) term order or via the Eigenvalue Method.
@@
    Factor the univariate polynomial.
For Ciminion and Hydra the designers conjectured that Step (2) dominates the complexity of a Gröbner basis attack. In addition, to estimate the complexity of this step it was assumed that the polynomial systems are regular or semi-regular [BFS04]. We on the other hand will construct DRL Gröbner bases for both designs via linear transformations and possibly a linear change of variables. Thus, we can directly proceed with Step (3). Typically, this step is also computationally cheaper than finding the DRL Gröbner basis for generic polynomial systems, which is the main reason why we yield better complexity estimations than the designers.
1.1 Our Results & Relationship to Other Works
The research goals of this work can be summarized in three bullet points:
  • Increase the understanding of polynomial models for MPC-friendly designs.
  • Provide clean cryptanalysis which avoids strong genericity assumptions.
  • Improve the complexity estimations of Ciminion and Hydra from a designer’s point of view.
Ciminion.
The PRF Ciminion is an iterated Feistel design based on the Toffoli gate [Tof80]. For the iterated Ciminion polynomial model we construct a DRL Gröbner basis via linear transformations, see Theorem 3.1. Hence, we trivialize the hardness assumption of Step (2).
It is worthwhile mentioning that in a recent work Bariant [Bar23] constructed the univariate LEX polynomial of Ciminion via iterated univariate polynomial multiplications. The complexities of constructing and factoring the LEX polynomial are linear-logarithmic in its degree. On the other hand, constructing the univariate LEX polynomial via Gröbner basis methods scales at least quadratic-logarithmic in its degree. Therefore, to the best of our knowledge Bariant’s attack currently constitutes the most competitive attack on certain Ciminion parameters. We also review this attack in more detail in Section 3.2.
There are two simple countermeasures against Bariant’s attack: increasing the number of rounds or additional key additions. For the latter, we also discuss in Section 3.2 how key additions can be performed while maintaining the Ciminion DRL Gröbner basis from Theorem 3.1. Additionally, in Section 3.3.1 we propose Ciminion2, a slightly modified variant which incorporates key additions that are compatible with Theorem 3.1.
In Section 3.4 we discuss Gröbner basis cryptanalysis of Ciminion and Ciminion2 respectively. Due to our DRL Gröbner basis we yield a clean analysis as well as state-of-the-art complexity estimations for Ciminion. In particular, we improve upon the designer’s original analysis. In Table 1 we present the minimal Ciminion round numbers that achieve $128$ bits of security for an ideal adversary [^1] In our setting an ideal adversary has access to a hypothetical matrix multiplication algorithm which achieves the linear algebra constant $\omega=2$. against the designer’s estimation (“Fully Substituted Model”), Bariant’s attack and our dedicated Eigenvalue Method for Ciminion and Ciminion2.
@@ 1.1 Our Results & Relationship to Other Works
Table 1: Minimal number of rounds required for Ciminion and Ciminion2 to achieve at least $128$ bits of security against various Gröbner basis attacks. All computations are performed over the prime $q=2^{127}+45$ with $\omega=2$.
Complexity (bits)
$r_{C}+r_{E}$ Bariant’s Attack [Bar23] Eigenvalue Method Fully Substituted Model [DGGK21, §4.4]
$33$ $46.67$ $63.09$ $130$
$66$ $81.22$ $129.09$ $262$
$112$ $128.47$ $221.09$ $446$
Hydra.
The heads of the Hydra can be considered as iterated Lai–Massey ciphers, where the key as well as the input are hidden from the adversary and only the output is released. In a recent work Steiner [Ste24] introduced the notion of polynomial systems in generic coordinates [CG21] to cryptographic polynomial models. In essence, if a polynomial system is in generic coordinates, then we have a proven complexity for the computation of the DRL Gröbner basis [CG21]. In [Ste24, §6.2] the iterated polynomial models of generalized Feistel ciphers were analyzed for being in generic coordinates. For many instances of Feistel ciphers generic coordinates can be verified via the rank of an associated linear system. In Theorem 4.1 we will construct such a linear system for the iterated Hydra model. Hence, with the aid of a computer algebra program like SageMath [Sag24] we yield a computer-aided proof for the complexity of DRL Gröbner basis computations. Additionally, if the linear system for generic coordinates verification has full rank, then we can perform a linear change of coordinates to transform the Hydra polynomial model into a DRL Gröbner basis, see Section 4.2.
Via the change of coordinates we produce a quadratic DRL Gröbner basis in $22\cdot r_{\mathcal{H}}-$ 2 polynomials and variables together with four additional quadratic polynomials. This leaves us with two possible routes to recover the key:
  1. (A)
    We can recompute the DRL Gröbner basis for $22\cdot r_{\mathcal{H}}+$ 2 quadratic polynomials.
  2. (B)
    We ignore the additional equations and just use the DRL Gröbner basis to construct a univariate polynomial.
The first approach was also considered by the Hydra designers [GØSW23, §7.4], although without the linear transformation to the DRL Gröbner basis. Under the semi-regularity assumption [BFS04] the designers derived the degree of regularity, which determines the complexity of the computation. In principle, we can apply the same strategy, but with one important benefit: After the linear transformation we only have to consider square-free monomials in a DRL Gröbner basis computation. This effectively saves both time and space. In analogy to Gröbner basis computations over $\mathbb{F}_{2}$ we call this the Boolean Semi-Regular Gröbner basis attack.
Alternatively, we can simply restrict to the DRL Gröbner basis and proceed with term order conversion via the probabilistic FGLM algorithm [FGHR14]. It turns out that this approach invalidates the designers’ claim that $r_{\mathcal{H}}=31$ rounds cannot be attacked by an ideal adversary below $128$ bits. In addition, via a dedicated Eigenvalue Method an ideal adversary can attack up to $r_{\mathcal{H}}=33$ rounds.
In Section 4.4 we discuss the complexity estimations for Hydra in more detail, a summary of the considered attacks is given in Table 2. Let $r_{\mathcal{H}}^{\ast}$ be the least round number which achieves $128$ bits of security against an ideal adversary. The Hydra designers require for the head round number that $r_{\mathcal{H}}=\left\lceil 1.25\cdot\max\left\{24,2+r_{\mathcal{H}}^{\ast}\right\}\right\rceil$ [GØSW23, §5.5]. In the original analysis the designers found that $r_{\mathcal{H}}^{\ast}=29$ and $r_{\mathcal{H}}=39$, however our analysis yields $r_{\mathcal{H}}^{\ast}=34$ which implies the increase $r_{\mathcal{H}}=45$ to maintain the originally intended security margin.
Table 2: Complexities of Gröbner basis attacks on Hydra which beat the designers’ analysis, and security margin reduction for full rounds. All computations are performed over the prime $q=2^{127}+45$ with $\omega=2$.
Complexity (Bits)
@@
$r_{\mathcal{H}}$ Term Order Conversion Eigenvalue Method Boolean Semi-Regular Estimate Semi-Regular Estimate [GØSW23]
$29$ $117.81$ $111.09$ $123.29$ $130.80$
$31$ $125.91$ $119.09$ $135.91$ $141.77$
$33$ $134.00$ $127.09$ $145.75$ $152.75$
$39$ $158.25$ $151.09$ $172.66$ $182.22$
Implementation.
We implemented the Ciminion and Hydra polynomial systems and their DRL Gröbner bases in the computer algebra systems SageMath [Sag24] and OSCAR [OSC24]. [^2] https://github.com/sca-research/Groebner-Basis-Cryptanalysis-of-Ciminion-and-Hydra.git In particular, for Hydra we have implemented the linear system for the generic coordinates verification as well as the linear change of coordinates to produce a quadratic DRL Gröbner basis. We stress that our SageMath implementations are able to handle practical parameters like $128$ bit prime numbers and full rounds Hydra with $r_{\mathcal{H}}=39$.
1.2 Organization of the Paper
In Section 2 we introduce the technical requirements of this paper. Ciminion and Hydra are formally introduced in Sections 2.1 and 2.2 respectively. In Section 2.3 we quickly review Gröbner basis algorithms and the notion of generic coordinates. As preparation for Hydra cryptanalysis, in Section 2.4 we consider polynomial systems $\mathcal{F}\subset K[x_{1},\dots,x_{n}]$ which contain a subset $\mathcal{F}_{\text{Bool}}=\{g_{1},\dots,g_{n}\}\subset\mathcal{F}$ with $\operatorname{LM}_{DRL}\left(g_{i}\right)=x_{i}^{2}$ for all $1\leq i\leq n$. For such polynomial systems we can reduce Gaussian elimination on the Macaulay matrix $M_{\leq d}$ to a submatrix $M_{\leq d}^{\text{Bool}}$ which only represents the square-free monomials. The complexity of square-free Gaussian elimination is derived in Theorem 2.10, and the one of the auxiliary division by remainder step is given in Corollary 2.12. Finally, in Section 2.5 we develop a dedicated strategy, called the Eigenvalue Method, to find a $\mathbb{F}_{q}$-valued key guess via the quadratic DRL Gröbner bases for Ciminion and Hydra.
In Section 3 we analyze the iterated Ciminion polynomial system. In particular, in Theorem 3.1 we construct a Ciminion DRL Gröbner basis via linear transformations. Bariant’s attack and compatibility of countermeasures with the DRL Gröbner basis are discussed in Section 3.2. Finally, in Section 3.4 we discuss Gröbner basis cryptanalysis of Ciminion.
In Section 4 we analyze the iterated polynomial system for the Hydra heads. A linear system to verify generic coordinates for Hydra is developed in Theorem 4.1. In case the linear system has full rank we discuss in Section 4.2 that a linear change of coordinates yields a quadratic DRL Gröbner basis together with four additional quadratic polynomials. Note that we do not perform this change of coordinates by hand, instead we outsource it to our SageMath or OSCAR implementations. In Section 4.4 we finish this paper with Gröbner basis cryptanalysis for Hydra.
@@ 2 Preliminaries
We denote fields with $K$ and their algebraic closure by $\overline{K}$. The finite field with $q$ elements is denoted as $\mathbb{F}_{q}$. If the field and the number of variables are clear from context, then we abbreviate the polynomial ring as $P=K[x_{1},\dots,x_{n}]$. Matrices $\mathbf{M}\in K^{m\times n}$ are denoted with upper bold letters and vectors $\mathbf{v}\in K^{n}$ by lower bold letters. Matrix-vector and matrix-matrix products are written as $\mathbf{M}\mathbf{v}$.
@@
Analogously, we denote the homogenization of ideals $I^{\text{hom}}=\left\{f^{\text{hom}}\mid f\in I\right\}$ and of finite systems of polynomials $\mathcal{F}^{\text{hom}}=\left\{f_{1}^{\text{hom}},\dots,f_{m}^{\text{hom}}\right\}$.
Let $f\in K[x_{1},\dots,x_{n}]$ be a polynomial, we can always decompose it as sum $f=f_{d}+\ldots+f_{0}$, where $f_{i}$ is homogeneous of degree $i$. We call $f_{d}$ the homogeneous highest degree component of $f$ and denote it as $f^{\text{top}}$. Analogously, we denote $\mathcal{F}^{\text{top}}=\left\{f_{1}^{\text{top}},\dots,f_{m}^{\text{top}}\right\}$. Note that we have $f^{\text{top}}=f^{\text{hom}}\mod\left(x_{0}\right)$.
Let $I\subset R$ be an ideal in a commutative ring, the radical of $I$ is defined as $\sqrt{I}=\left\{f\in R\mid\exists n\in\mathbb{Z}_{\geq 1}\!:f^{n}\in I\right\}$. It is well-known that the radical is also an ideal.